# Generating synthetic data for neural operators

### v1

@@ 5 Conclusion and future work
Using deep learning to solve PDEs has been very promising in recent years. Here, we propose a method that eliminates the need to repeatedly solve a PDE for obtaining training data used in neural operators by first generating the unknown solution and then computing the right-hand side of the equation. Although we exclusively provide theoretical motivation and numerical experiments for second-order elliptic PDEs, this concept can be extended to other types of PDEs where the solution space is known beforehand, enabling the construction of representative functions for such solution spaces. This method could open up the possibility of obtaining good predictions for PDE problems using data-driven neural operators, for which the training data does not require classical numerical solvers to generate. We stress that our synthetic data generation approach is computationally efficient, particularly compared to solving new PDE problems numerically to generate training data for each new problem instance. We believe that our approach is important towards reaching the ultimate goal of using deep learning to solve PDEs that are outside the scope of what we can solve using classical numerical solvers. We note that as a by product, our method eliminates sources of error coming from numerically solving PDE problems; instead, our synthetic training data is of the form of exact solutions to a problem on a fixed-size grid.

### v3

@@ 5 Limitations, conclusion and future work
Conclusion and future work. Using deep learning to solve PDEs has been very promising in recent years. Here, we propose a method that in some settings could eliminate the need to repeatedly solve a PDE for obtaining training data used in training neural operators by first generating the unknown solution and then computing the right-hand side of the equation. Although we exclusively provide theoretical motivation and numerical experiments for second-order elliptic PDEs, this concept could be extended to other types of PDEs where the solution space is known beforehand, enabling the construction of representative functions for such solution spaces. This method could open up the possibility of obtaining good predictions for PDE problems using data-driven neural operators, for which the training data does not require classical numerical solvers to generate. We stress that our synthetic data generation approach is computationally efficient, particularly compared to solving new PDE problems numerically to generate training data for each new problem instance. We believe that our approach is an important step towards reaching the ultimate goal of using deep learning to solve PDEs that are intractable using classical numerical solvers. We also note that as a by-product, our method eliminates sources of error coming from numerically solving PDE problems; instead, our synthetic training data is of the form of exact solutions to a problem on a fixed-size grid. As future work, a promising direction is to use our synthetic data for pre-training, followed by fine-tuning on a few task-specific examples, in line with recent transfer learning approaches explored for elliptic problems (e.g. [10] and [30]).